40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 22 17:17:30 2013, annalisa, Update, endtable upgrade, End table picture YendTable.jpg
    Reply  Tue Jul 23 00:51:46 2013, Koji, Update, endtable upgrade, End table picture P7224222.JPGP7224223.JPGP7224224.JPGP7224238.JPG
       Reply  Tue Jul 23 19:37:08 2013, Annalisa, Update, endtable upgrade, End table picture IMG950073.jpgIMG950072.jpg
    Reply  Wed Sep 25 19:34:51 2013, rana, Update, SUS, Problems with ETMY Optical Lever ETMY-BadHeNe.pdf
       Reply  Thu Sep 26 11:00:51 2013, Steve, Update, SUS, Problems with ETMY Optical Lever 
          Reply  Thu Sep 26 21:55:08 2013, rana, Update, SUS, Problems with ETMY Optical Lever 
Message ID: 9160     Entry time: Wed Sep 25 19:34:51 2013     In reply to: 8892     Reply to this: 9165
Author: rana 
Type: Update 
Category: SUS 
Subject: Problems with ETMY Optical Lever 

I went down to investigate the issue with the extra noise that I found in the ETMY optical lever yesterday. There were several problems with the optical layout down there - I'm not sure if I remember them all now.

  1. Beam reflected from OL QPD not dumped.
  2. OL QPD set normal to the steering mirror so that the back reflection goes into the vacuum chamber.
  3. HeNe laser mount only dogged with 2 dogs. Needs 3. Looks like some said "Aw, that's not goin' nowhere. Let's just leave that there pard!"
  4. First lens downstream of the laser had 2 screws and washers, but neither was even finger tight! They were loose by more than 1 full turn.
  5. Second lens was clipping. Beam was so far off center that this lens was being used to steer the beam by a few inches on the QPD.
  6. Extra reflections from ingoing beam (I don't know which surfaces) randomly landing on green & red optics.
  7. Lenses for the HeNe mode matching are coated for 1064 nm. HeNe is 633 nm, so these lenses must be replaced to reduce the reflections.

The main noise issue, however, appears to be not a layout issue at all. Instead its that the laser intensity noise has gone through the roof. See attached spectra of the quadrants (this is the way to diagnose this issue).

I'll ask Steve to either heal this laser or swap it out tomorrow. After that's resolved we'll need another round of layout fixing. I've done a couple of hours today, but if we want a less useless and noisy servo we'll have to do better.

NOTE: by looking at the OL quadrants, I've found a noisy laser, but this still doesn't explain the excess noise in the ETMX. That was the one that has a noisier error signal, not ETMY. By the coherence in the DTT, you can see that the ETMY OL is correctly subtracting and normalizing out the intensity noise of the laser. Seems like the ETMX electronics might be the culprit down there.

Attachment 1: ETMY-BadHeNe.pdf  171 kB  Uploaded Wed Sep 25 20:50:17 2013  | Hide | Hide all
ETMY-BadHeNe.pdf
ELOG V3.1.3-